OpenStudy (anonymous):

Solve the equation for w: wxy + z + 1 = 6

OpenStudy (jiteshmeghwal9):

Subtract z & 1 both sides\[wxy+z+1-z-1=6-z-1\]\[wxy=6-z-1\]Now divide \(xy\) both sides\[\dfrac{wxy}{xy}=\dfrac{6-z-1}{xy}\]

OpenStudy (anonymous):

thanks got it!

OpenStudy (jiteshmeghwal9):

wc :)

OpenStudy (anonymous):

can you help me out with one more like that?